OpenStudy (anonymous):

Solve : 50^2 + 60^2 = N * sq.rootsign625 Help please, thx in advanced!

OpenStudy (anonymous):

50^2+60^2=25N 2500+3600=25N 6100=25N 6100/25=N 244=N N=244